# Robust surface superconductivity and vortex lattice in the Weyl semimetal -PtBi2

### v1

@@ Abstract
The layered compound -PtBi2 is a topological semimetal with Fermi arcs at the surface joining bulk Weyl points. Recent work has found a peculiar surface superconducting state with a critical temperature two orders of magnitude larger than the bulk value. However, no superconducting vortices have been identified, raising questions about the robustness of the phase coherence in this new surface superconducting state. Here we measure the superconducting gap and visualize the vortex lattice through very low temperature Scanning Tunneling Microscopy (STM). We find vortices that are anomalously mobile and show that the extremely weak electrostatic interaction between the vortex charge and the STM tip is sufficient to drag the vortices of the surface superconductor. Our results show that surface superconductivity with K and T is extraordinarily robust in -PtBi2. Furthermore, quasiparticle scattering is enhanced at the position in the reciprocal space of the Fermi arcs, suggesting a connection between Weyl points and surface superconductivity.
@@
Many semimetals present Weyl or Dirac nodes in their electronic band structures [1, 2]. The bulk-boundary correspondence often dictates the presence of topologically protected bound states at the surface. This is particularly interesting when it occurs with electron-hole symmetric quasiparticles in a superconductor, because it allows to engineer systems providing Majorana modes [3]. However, few of these semimetals are superconductors in the bulk. This has lead to proposals to study heterostructures consisting of a thin layer of a topological semimetal on top of a bulk superconductor [4, 5, 2, 6, 7, 8]. Superconductivity is then induced in the surface state of the semimetal by proximity. Contrasting these proposals, it could also happen that superconductivity is intrinsically realized on the very same topologically protected surface state of a bulk material with Dirac or Weyl nodes [9, 10, 11]. Recent experiments and calculations suggest that this situation could be realized in the layered Weyl semimetal PtBi2 [12, 13, 14, 15, 16, 17].
@@
PtBi2 crystallizes in a layered trigonal structure without inversion symmetry [18]. The band structure contains wide bands crossing the Fermi level at the upper and lower borders of the Brillouin zone. However, the center of the Brillouin zone is gapped. This provides a semimetallic character showing peculiar transport properties as a huge and non-saturating magnetoresistance [19]. The electronic band structure contains Weyl points at approximately 50 meV above the Fermi level [1, 2, 19]. The calculated electronic band structure agrees well with the angle resolved photoemission (ARPES), quasiparticle scattering (QPI) and quantum oscillations experiments [17, 12, 20, 21]. Furthermore, ARPES finds Fermi arcs in surface bands located close to the projection of the Weyl points to the surface plane [12, 20]. At the same time, indications for surface superconductivity with critical temperatures well above 10 K, strongly enhanced with respect to the bulk ( K) [22, 18, 23] and presenting superconducting gap values of several tens of meV have been proposed based on ARPES and Scanning Tunneling Microscopy (STM) [12, 13, 14, 15, 16, 17]. However, the features in the density of states observed with STM provide widely different values of the superconducting gap and recent ARPES measurements do not show any signature of superconducting gap opening on the Fermi arcs within an energy range down to below 1 meV [15, 20]. What is more, the observed gaps close to the Fermi level persist up to very strong magnetic fields (up to 12 T) and there are no reported observations of a vortex lattice nor of the Josephson effect [15, 24]. To better understand the eventual connection of superconductivity and Fermi arcs, it is important to establish the basic properties of surface superconductivity in PtBi2.
@@
We use here dilution refrigerator STM under magnetic fields to study PtBi2. We find homogeneous superconductivity down to atomic scale with a critical temperature of 2.9 K and the gap size and temperature dependence following BCS theory. We observe superconducting vortices and study the behavior of the vortex lattice as a function of magnetic field and temperature, finding an anomalously high vortex mobility which is due to the exotic two-dimensional character of superconductivity. QPI shows features of the Fermi arcs at the superconducting quasiparticle peaks.
@@
High-quality single crystals of PtBi2 were grown using a Bi-rich flux [20], as described in the Supplementary Material [25]. When probed macroscopically either by resistivity or by magnetization, there are no hints of bulk superconductivity down to K. The residual resistance ratio is of RRR, showing that the crystals have minimal disorder scattering. We used a dilution-refrigerator-STM with an energy resolution below 8 eV described in Refs. [26, 27], and the data treatment software described in Refs. [28, 29, 30]. Further details on the STM data acquisition are described in the Supplementary Material [25]. The cleaving plane is located in between two Bi layers, as shown in Fig. 1 (a). The two possible cleaving surfaces present a trigonal arrangement of Bi atoms, one in which Bi atoms located at two nonequivalent sites (Bi1 and Bi2, termination A) form intertwined corrugated hexagons, and another where the Bi are located on a single plane (Bi3, termination B). The resulting STM topographic images show hexagonal patterns (Fig. 1(b,c)), characteristic of each termination. Both terminations provide the same superconducting and vortex lattice properties, described below.
@@
In Fig. 1(e) we show a typical current vs voltage curve for -PtBi2 at 0.1 K obtained on atomically flat areas. We note that the current is linear as a function of the voltage, suggesting excellent metallic behavior. We also note the clear and small superconducting gap that opens and a zero current inside the gap. To visualize this better, we show the tunneling conductance vs bias voltage in Fig. 1(f). We find large quasiparticle peaks and zero density of states inside the gap. We can fit the tunneling conductance using a modified BCS density of states, with Re, convoluted with the derivative of the Fermi function at 0.1 K. For a perfectly isotropic s-wave BCS superconductor with gap , when and zero elsewhere. Quite often, however, the superconducting gap is not perfectly homogeneous and there is a certain distribution of gap values over the Fermi surface [31, 32, 33, 34, 35, 36]. We find here a Gaussian distribution centered around meV with a width of 0.07 meV.
@@
The spatial dependence of the superconducting tunneling conductance at atomic scale is shown in Fig. 2(a-c). The topography shown in Fig. 2(a) provides a typical atomically resolved STM image, containing a few defects, which appear as atomic size adatoms or interstitials (white spots) and as atomic vacancies (black spots). In Fig. 2(b) we represent a line scan of the height measured by STM. We see that the atomic corrugation is small, of a few pm. At the atomic size defects, there are peaks or troughs in the surface of a few tens of pm. Along the whole path, the superconducting gap remains of the same size (Fig. 2(c)). The defects do not produce visible states at low energies well inside the superconducting gap.
@@
The temperature dependence of the superconducting tunneling conductance is shown in Fig. 2(d,e). We can fit the whole temperature dependence using the discussed previously, convoluted by the derivative of the Fermi function at the temperature of the measurement, without any adjustment parameters up to about 1 K. Above that temperature, the value of the peak is reduced in , which gives the provided in the right panel of Fig. 2(d). We plot the peak position in as a function of temperature in Fig. 2(e) and find an excellent agreement with BCS theory (red line) up to the critical temperature of T K. With this value of Tc we can estimate the zero temperature superconducting gap using BCS theory, and we find meV, which coincides with better than within 10%.
We present the behavior of superconducting vortices, obtained by mapping the zero bias conductance as a function of the position, for different values of the magnetic field and at 0.1 K in Fig. 3. In Fig. 3(a) we show an isolated vortex. The tunneling conductance outside the vortex is close to its zero field value, and at the core of the vortex we observe a flat tunneling conductance (Fig. 3(b)).
In Fig. 3(c) we show a zero bias tunneling conductance map obtained at 0.4 T which is representative for the typical vortex lattice behavior we observe. In Fig. 3(d) we show the corresponding topography, we can identify atomically flat terraces on the upper part of the topography and nanometer-sized flakes which surge on top of the atomically flat terraces on the lower part. Concurrently, on the upper part of the zero bias tunneling conductance map (Fig. 3(c)) we find a homogeneous superconducting tunneling conductance, with curves similar to those found at zero field. On the lower part of Fig. 3(c) we observe, by contrast, the vortex lattice. The tunneling conductance curves in and outside vortices are similar to those shown in Fig. 3(a,b). We note that we cannot distinguish between the tunneling conductance outside vortex cores on nanometer-sized flakes and on the atomically flat areas. This behavior is found all over the surface of -PtBi2 in hundreds of fields of view, in different samples, and for different surface terminations on the atomically flat terraces—atomically flat terraces present no trace of vortices, whereas nanometer-sized flakes provide a neat vortex lattice.
We show in Fig. 3(e,f) a zero bias tunneling conductance map on a different field of view and for two values of the magnetic field. We see that the vortex density increases with the magnetic field. The intervortex distance as a function of the magnetic field (Fig. 3(g)) follows Abrikosov behavior for a hexagonal vortex lattice [37, 38]. Indeed, although the vortex lattice shows some degree of disorder, the Fourier transform mostly presents a hexagonal pattern (insets of Fig. 3(e,f)). By extrapolating the zero bias tunneling conductance outside vortices to high fields, we estimate an upper critical field 1.8 T, applied perpendicular to the surface, which provides an in-plane coherence length of about nm.
From our result, we see that vortices must be found all over the sample surface, including atomically flat terraces, because the very large magnetic fields we apply here inherently create vortices in a superconductor. The apparent absence of vortices in STM on flat terraces should be due to peculiar vortex pinning properties.
To better understand this point, let us consider the electrostatic interaction of vortices with a non-magnetic tip. As shown in Ref. [39], there is a redistribution of charge due to the differences in the chemical potential between superconducting and normal quasiparticles. Metallic screening does not fully eliminate the charge. This results in an electric dipole oriented perpendicular to the surface whose strength is of the order of where is the elementary charge and Å(which gives D, one Debye, see Fig. 11 in the Supplementary Material). Such a minute electric dipole has been often considered previously and has an influence in several properties of superconductors, such as the Hall effect in the mixed state [40, 41, 42, 43, 44]. To estimate the force generated by the electric field of the tip on the vortex, we first consider the corresponding interaction energy, . The dipole sits on the vortex core, but its field extends over a distance of order of the superconducting coherence length . We now consider a tip which is at a distance smaller than and is blunt and large in size (see for a schematic picture Fig. 11 in the Supplementary Material). The effective electric field is then , where is the voltage applied to the tip. We can then estimate the force acting on the vortex as . For V, nm we obtain fN.
The pinning force can be roughly estimated by considering the free energy per unit volume from the thermodynamic critical field and the vortex volume [45]. Estimates are often of order of N/m [45, 46]. In PtBi2, the bulk is normal, which implies that superconductivity is restricted to a small layer close to the surface. Taking a thickness for the superconducting layer of 1 nm, we estimate fN (see Fig. 11 in the Supplementary Material), which is one order of magnitude below our estimate of . If superconductivity is restricted to a the thin layer occupied by a surface band, the thickness should be similar or smaller than a unit cell size, which is of 0.617 nm along the c-axis.
On the other hand, on nanometer-sized flakes, the vortex length increases, which also increases the pinning force, up to .
Thus, we conclude that on atomically flat surfaces, and vortices move together with the tip and remain undetected (further details provided in the Supplementary Material). However, when , the vortex lattice remains pinned and does not move during scanning. Then, STM reveals the presence of vortices.
Such a highly mobile vortex lattice should be very sensitive to temperature. Indeed, increasing temperature to about 2 K, well below , leads to a completely blurred vortex configuration (see Supplementary Material), similar to the features observed with STM in vortex liquid phases [47, 48, 49, 50, 51].
It is interesting to remark that the anisotropy of the upper critical field measured in thin flakes of PtBi2 suggests that the superconducting coherence length is twenty times smaller along the c-axis than in plane, [22]. Superconductivity is thus indeed restricted to a surface layer and is possibly even smaller than estimates given above. Furthermore, the coherence length perpendicular to the surface is so small, that the in-gap density of states induced by proximity from the normal bulk is practically absent in the superconducting density of states at the surface.
Thus, we have established that robust superconductivity with K exists on the surface of -PtBi2. STM provides a spatially continuous superconducting gap with a temperature dependence following BCS theory. In addition, we show that the Cooper pair wave function is uniquely valued, through the observation of flux quantization on the vortex lattice. Finally, as shown in the Supplementary Material, we observe the Josephson effect when using a superconducting tip.
To investigate the relationship between superconductivity and the Fermi arcs found in ARPES [12, 20], we have performed QPI maps of the tunneling conductance on an atomically flat terrace with sufficient atomic size defects to obtain scattering (on the area shown in Fig. 2(a)). As shown previously, the tunneling conductance vanishes within the gap, so that the tunneling conductance maps inside the gap are essentially zero. However, exactly at the quasiparticle peaks position, we find clear signatures of QPI. This suggests that the superconducting density of states varies spatially through quasiparticle scattering. This variation shows changes in the reciprocal space gap structure induced by impurity scattering. We show the Fourier transform of the tunneling conductance map at in Fig. 4(a). A similar result is found for . We find several scattering wave vectors agreeing with previous QPI and ARPES experiments at energies above the superconducting gap edge [17, 20]. By comparing with DFT (Fig. 4(b,c)) we find, at the superconducting gap edge, scattering wave vectors (, , and ) which coincide with wave vectors joining Fermi arcs that connect the projections of Weyl points to the surface.
@@
Whereas we observe superconductivity for temperatures below 3 K, compatible with the recent ARPES experiments of Ref. [20], the temperature range for earlier claims of superconductivity in -PtBi2 is reported to be even far above the Tc we observe here [12, 13, 14, 15, 16]. The repeated observation of atomically flat areas in our experiment, always presenting the same gap values, shows that the superconducting properties observed here must be due to intrinsic two-dimensional superconductivity at the surface. It remains unclear what is causing the surface superconductivity. It could be small shifts in the band structure leading to enhanced density of states and/or changes in the phonon structure close to the surface [9, 10]. Interestingly, monolayer PtBi2 has been proposed to be a ferroelectric metal due to phonon softening [52]. Furthermore, it has been recently proposed that normal metals with Weyl points can show phonon mediated surface superconductivity [11]. As we have now shown that such two-dimensional superconducting state exists and is robust, it is also interesting to ask about the relationship with the topological properties of the electronic band structure.
@@
Our QPI measurements show that the superconducting gap opens at the Fermi arcs. For superconducting Fermi arcs that are projections of bulk Weyl points, Majorana modes are expected inside vortices. These should be visible by a peak in the tunneling conductance inside the vortex core [3, 4, 5, 2, 6, 7, 8]. But even in absence of Majorana modes, superconductors in the clean limit present pronounced peaks in the tunneling conductance inside vortex cores [53, 54, 34, 50, 55, 56, 57]. Their absence in Fig. 3(b) suggests that the electronic mean free path on the nanometer-sized flakes is of the order or slightly smaller than the superconducting coherence length, nm. This is considerably smaller than the bulk mean free path, which is of 200 nm or larger, as estimated from the residual resistance ratio of our samples. Thus, the two-dimensional electronic band structure at the surface of the nanometer-sized flakes is sufficiently well developed locally to carry surface superconductivity, but the mean free path is not sufficiently large to provide vortex cores with localized states inside.
Vortices on atomically flat terraces, where we also took the QPI data showing the connection between superconductivity and Fermi arcs (Fig. 4(a)), are probably in the clean limit and carry bound states, which could realize Majorana modes [11]. The charge of vortices carrying Majorana modes has been investigated theoretically. It is found that, depending on the different possible chiral p-wave superconducting states, vortices carrying Majorana fermions might be uncharged or charged [58, 59]. The latter are expected to present a considerably enhanced charge with respect to conventional s-wave superconductors [58, 59]. Our data suggest that the vortices on atomically flat terraces carry a substantial charge and have therefore a strong interaction with the tip.
In summary, we have found robust surface superconductivity in -PtBi2 with a Tc much larger than the bulk and which is connected to the Fermi arcs joining the projection of bulk Weyl points. The two-dimensional superconductivity confers exceptional vortex mobility, an aspect which should be taken into account to address unconventional superconductivity in topological edge states and will be of help to design experiments on interaction among unconventional bound states in superconductors.

### v2

# Robust two-dimensional surface superconductivity and vortex lattice in the Weyl semimetal -PtBi2

@@ Abstract
The layered compound -PtBi2 is a topological semimetal with Fermi arcs at the surface joining bulk Weyl points. Recent work has found signatures of surface superconductivity consisting of gap openings compatible with a critical temperature orders of magnitude larger than the bulk value. However, no superconducting vortices have been identified, raising questions about the robustness of the phase coherence. Here we use very low temperature Scanning Tunneling Microscopy (STM) and find robust superconductivity with T2.9 K and H1.8 T linked to the Fermi arcs. We observe quantized superconducting vortices, demonstrating two-dimensional macroscopic quantum phase coherence.
@@
Many semimetals present Weyl or Dirac nodes in their electronic band structures [1, 2]. The bulk-boundary correspondence often dictates the presence of topologically protected bound states at the surface, which are of particular interest in superconductors because of possible non-Abelian modes [3]. Nevertheless, few of these semimetals are superconductors in the bulk. This has lead to proposals to study heterostructures consisting of a thin layer of a topological semimetal on top of a bulk superconductor [4, 5, 2, 6, 7, 8]. Superconductivity is then induced in the surface state of the semimetal by proximity. Contrasting these proposals, it could also happen that superconductivity is intrinsically realized on the very same topologically protected surface state of a bulk material with Dirac or Weyl nodes [9, 10, 11]. Recent experiments and calculations suggest that this situation could be realized in the layered Weyl semimetal PtBi2 [12, 13, 14, 15, 16, 17, 18]. The realization of such surface two-dimensional superconductivity with topologically protected states could have important applications in quantum devices [19, 20, 21] and complement efforts to understand two-dimensional superconductivity in rhombohedral, twisted bilayer and trilayer graphene [22, 23, 24], few-layer single crystals [25, 26] and ultra-thin films [27, 28, 29, 30].
@@
PtBi2 crystallizes in a layered trigonal structure without inversion symmetry [31]. The band structure contains wide bands crossing the Fermi level at the upper and lower borders of the Brillouin zone. However, the center of the Brillouin zone is gapped. This provides a semimetallic character presenting a huge and non-saturating magnetoresistance [32]. The electronic band structure contains Weyl points at approximately 50 meV above the Fermi level [1, 2, 32]. The calculated electronic band structure agrees well with the angle resolved photoemission (ARPES), quasiparticle scattering (QPI) and quantum oscillations experiments [18, 13, 33, 34]. Furthermore, ARPES finds Fermi arcs in surface bands located close to the projection of the Weyl points to the surface plane [13, 33]. Specific heat of bulk PtBi2 does not present traces of superconductivity down to 0.5 K. There are large resistive transitions with well below 1 K in thin flakes of PtBi2 obtained by exfoliation [16, 35, 14, 31, 36]. Other indications for surface superconductivity present ’s well above 10 K, associated with superconducting gap values of several tens of meV and are based on ARPES and Scanning Tunneling Microscopy (STM) measurements [13, 15, 16, 17, 18]. The values of the superconducting gap obtained from ARPES and STM vary widely among different experiments, with the most recent ARPES measurements showing no signature of a superconducting gap opening on the Fermi arcs at temperatures of 3 K and within an energy range down to below 1 meV [16, 33]. What is more, the gaps observed around the Fermi level in STM often persist up to very strong magnetic fields (up to 12 T) and there are no reported observations of a vortex lattice nor of the Josephson effect [16, 37]. To better understand the eventual connection of superconductivity and Fermi arcs, it is important to demonstrate the presence of vortices and thus flux quantization in the surface superconductivity of PtBi2.
@@
Here, we use dilution refrigerator STM under magnetic fields to study PtBi2. We find homogeneous superconductivity down to atomic scale with a critical temperature of 2.9 K and the gap size and temperature dependence following BCS theory. We observe superconducting vortices and study the behavior of the vortex lattice as a function of magnetic field and temperature, finding an anomalously high vortex mobility which is due to the two-dimensional character of superconductivity. QPI shows features of the Fermi arcs at the superconducting quasiparticle peaks.
@@
High-quality single crystals of PtBi2 were grown using a Bi-rich flux [33], as described in the Supplemental Material [38]. When probed macroscopically either by resistivity or by magnetization, there are no hints of bulk superconductivity down to K (see Fig. 1(d) and Supplemental Material [38]). The residual resistance ratio is of RRR, showing that the crystals have minimal disorder scattering. We used a dilution-refrigerator-STM with an energy resolution below 8 eV cooling to 0.1 K, described in Refs. [39, 40], and the data treatment software described in Refs. [41, 42, 43]. Further details on the STM data acquisition are described in the Supplemental Material [38]. The cleaving plane is located in between two Bi layers, as shown in Fig. 1(a). The two possible cleaving surfaces present a trigonal arrangement of Bi atoms, one in which Bi atoms located at two nonequivalent sites (Bi1 and Bi2, termination A) form intertwined corrugated hexagons, and another where the Bi are located on a single plane (Bi3, termination B). The resulting STM topographic images show hexagonal patterns (Fig. 1(b,c)), characteristic of each termination, as found previously in Ref. [16]. Both terminations provide the same superconducting and vortex lattice properties described below.
@@
In Fig. 1(e) we show a characteristic current vs voltage curve for -PtBi2 at 0.1 K obtained on atomically flat areas. We note that the current is linear as a function of the voltage, suggesting excellent metallic behavior. We also note the clear and small superconducting gap that opens and a zero current inside the gap. To visualize this better, we show the tunneling conductance vs bias voltage in Fig. 1(f). We find large quasiparticle peaks and zero density of states inside the gap. We fit the tunneling conductance using a modified BCS density of states, with Re, convoluted with the derivative of the Fermi function at 0.1 K. For a perfectly isotropic s-wave BCS superconductor with gap , when and zero elsewhere. In systems with an anisotropic band structure, however, the superconducting gap is often not perfectly homogeneous and there is a certain distribution of gap values over the Fermi surface [44, 45, 46, 47, 48, 49]. We find here a Gaussian distribution centered around meV with a width of 0.07 meV.
@@
The spatial dependence of the superconducting tunneling conductance at atomic scale is shown in Fig. 2(a-c). The topography shown in Fig. 2(a) provides an atomically flat surface, containing a few defects, which appear as atomic size adatoms or interstitials (white spots) and as atomic vacancies (black spots). In Fig. 2(b) we represent a line scan of the height measured by STM. We see that the atomic corrugation is small, of a few pm. At the atomic size defects, there are peaks or troughs in the surface of a few tens of pm. Along the whole path, the superconducting gap slightly fluctuates (Fig. 2(c)). The defects do not strongly influence the tunneling conductance.
@@
The temperature dependence of the superconducting tunneling conductance is shown in Fig. 2(d,e). We can fit the whole temperature dependence using the discussed previously, convoluted by the derivative of the Fermi function at the temperature of the measurement, without any adjustment parameters up to about 1 K. Above that temperature, the value of the peak is reduced in , which gives the provided in the right panel of Fig. 2(d). We plot the peak position in as a function of temperature in Fig. 2(e) and find an excellent agreement with BCS theory (red line), obtained by solving the self-consistent gap equation with T K and meV. The latter coincides with better than within 10%.
We present the superconducting vortex lattice, obtained by mapping the zero bias conductance as a function of the position, for different values of the magnetic field and at 0.1 K in Fig. 3. In Fig. 3(a) we show a zero bias conductance map with a single vortex (for more details see Supplemental Material [38]). In Fig. 3(b,c,d) we show zero bias tunneling conductance maps obtained at 0.4 T, 0.8 T and 1.3 T respectively. The intervortex distance as a function of the magnetic field (Fig. 3(e)) follows Abrikosov behavior for a hexagonal vortex lattice [50, 51]. The Fourier transforms of vortex lattice images mostly present a hexagonal patterns, although the vortex lattice is disordered (insets of Fig. 3(c,d)). As we discuss in the Supplemental Material [38], we measure the upper critical field using the magnetic field dependence of the tunneling conductance in between vortices and the magnetic field dependence of the radial behavior of the tunneling conductance around vortices, finding 1.8 T and an in-plane coherence length of about nm.
Interestingly, we find extended areas where we do not observe vortices under magnetic fields, see left part of Fig. 3(a) and portions of Fig. 3(b,c,d). Previous work focused on such areas and did not find indications for a vortex lattice[16, 18, 52]. In Fig. 3(f) we show the topography of the field of view shown in Fig. 3(b), see Supplemental Material for topographies in other fields of view. The areas where we do not observe vortices are atomically flat, and the areas where we observe vortices present a surface with a corrugation of fractions of a nm. The superconducting tunneling conductance is equivalent in both areas. We show the tunneling conductance far from a single isolated vortex in both areas in Fig. 3(g) (red and orange circles). Furthermore, the superconducting tunneling conductance averaged over a single unit cell of the vortex lattice in the areas where we observe vortices is equivalent to the tunneling conductance found in the areas where we do not observe vortices (Fig. 3(h)). We provide further line scans in the Supplemental Material. All this, together with scanning electron microscopy structural and elemental characterization provided in Supplemental Material [38], show that the areas with an irregular topography are nanometer-sized flakes of the same sample produced during the cleave.
Vortices are also present on atomically flat terraces, as demonstrated by the equivalent tunneling conductance in both areas at high magnetic fields (Fig. 3(h)). The absence of individual vortex resolution in the tunneling conductance maps on flat terraces can be due to enhanced vortex mobility in those areas. An electrostatic interaction between the STM tip and vortex-induced charge [53] can generate a force exceeding the weak pinning force of vortices in a two-dimensional superconducting surface layer, causing vortices to move below the tip. In contrast, on thicker nanometer-sized flakes, there is enhanced pinning due to an increased effective vortex length, stabilizing the vortex lattice and allowing for its observation in tunneling conductance maps. We provide more details about vortex pinning in the Supplemental Material [38], including conductance maps showing thickness-dependent vortex pinning and tip-induced vortex motion.
Thus, we observe the main characteristic aspects of a two-dimensional superconducting vortex lattice. First, vortices are highly mobile. Second, as we show in the Supplemental Material [38], the vortex lattice melts when increasing temperature well below Tc, as found previously for two dimensional vortex lattices [54, 55, 56, 57, 58]. Third, the magnetic field penetration is very large due to the expansion of the magnetic profile of the two-dimensional vortex [59, 51, 50] and we indeed find little or no dependence of the vortex density from the magnetic field history.
As we discuss above and in the Supplemental material [38], there is no signature of superconductivity in bulk measurements in the field and temperature range discussed in Figs. 2,3. The anisotropy of the upper critical field () [35] shows that superconductivity in PtBi2 is confined to a thin surface layer. The very short coherence length along the c-axis suppresses proximity effects from the normal bulk, leaving the surface density of states essentially unaffected. Therefore, both the superconducting density of states and the vortex lattice are a two-dimensional surface property. We have thus established that robust two-dimensional superconductivity with K exists on the surface of -PtBi2. The Cooper pair wave function is uniquely valued, as shown by the observation of flux quantization and the vortex lattice. Furthermore, as shown in the Supplemental Material [38], we observe the Josephson effect when using a superconducting tip.
@@
Whereas we observe superconductivity for temperatures below 3 K, compatible with the recent ARPES experiments of Ref. [33] and the STM work of Ref. [52], the temperature range for earlier claims of superconductivity in -PtBi2 is reported to be even far above the Tc we observe here [13, 14, 15, 16, 17]. The repeated observation of the same gap values in our experiment shows that the superconducting properties observed here are reproducible. As we have now shown that such two-dimensional superconducting state exists and is robust, it is also interesting to ask about the relationship with the topological properties of the electronic band structure.
To investigate the relationship between surface two-dimensional superconductivity and the Fermi arcs found in ARPES [13, 33], we have performed QPI maps of the tunneling conductance voltage on an atomically flat terrace with sufficient atomic size defects to obtain scattering (in the area shown in Fig. 2(a)). As shown previously, the tunneling conductance vanishes within the gap, so that the tunneling conductance maps inside the gap are essentially zero. However, exactly at the quasiparticle peaks position, we find clear signatures of QPI. This suggests that the superconducting density of states varies spatially through quasiparticle scattering. This variation shows changes in the measured gap structure induced by impurity scattering. We show the Fourier transform of the tunneling conductance map at in Fig. 4(a). A similar result is found for . We find several scattering wave vectors agreeing with previous QPI and ARPES experiments at energies above the superconducting gap edge [18, 33]. By comparing with Density Functional Calculations (DFT, Fig. 4(b,c)) we find, at the superconducting gap edge, scattering wave vectors (, , and ) which coincide with wave vectors joining Fermi arcs that connect the projections of Weyl points to the surface.
@@
These QPI measurements show that the superconducting gap opens at the Fermi arcs, in agreement with theoretical predictions [60]. For superconducting Fermi arcs that are projections of bulk Weyl points, Majorana modes are expected inside vortices. These should be visible by a peak in the tunneling conductance inside the vortex core [3, 4, 5, 2, 6, 7, 8]. But even in absence of Majorana modes, superconductors in the clean limit present a large density of states inside vortex cores [61, 62, 47, 57, 63, 64, 65]. We do find a flat tunneling conductance inside vortex cores which suggests that the electronic mean free path on nanometer-sized flakes is comparable to or smaller than the coherence length (13 nm), smearing localized vortex core states despite surface superconductivity being present. In contrast, vortices on atomically flat surfaces are most likely in the clean limit and may host bound states, potentially including Majorana modes [11]. The charge of vortices carrying Majorana modes has been investigated theoretically. Depending on the particular superconducting state, vortices carrying Majorana fermions are expected to present a considerably enhanced charge with respect to conventional s-wave superconductors [66, 67]. Our data suggest that the vortices on atomically flat terraces carry a substantial charge and have therefore a strong interaction with the tip.
In summary, we have found robust surface superconductivity in -PtBi2 connected to the Fermi arcs joining the projection of bulk Weyl points. An interesting prospect is to better understand the pairing mechanism leading to two-dimensional surface superconductivity. Changes in the phonon structure close to the surface could play a role [9, 10]. Interestingly, monolayer PtBi2 has been proposed to be a ferroelectric metal due to phonon softening [68]. Furthermore, it has been recently proposed that metals with Weyl points can show topological surface superconductivity [11]. Superconductivity could be favored by the enhanced electronic density of surface states, in a distinct but somewhat comparable situation to the flat band surface state in rhombohedral graphene [23, 69, 70]. -PtBi2 provides an exceptional case of two-dimensional superconductivity arising at the surface state of a topological semimetal.
Note: During the final stages of the preparation of this work, we have become aware of an independent similar work which reports the observation of bulk superconductivity with 0.35 K and surface superconductivity with 3 K. Only atomically flat surfaces are found, without an observable vortex lattice, possibly as a consequence of different cleaving methods [52].